# Range Retrieval with Graph-Based Indices

### v2

@@ 4. Algorithms for Range Search on Graph-Based ANNS Indices
In this section we introduce our range search algorithms. We first introduce the data structure and baseline algorithm, a standard beam search. Since the standard beam search is well known to produce excellent results for top- queries, some range queries (those with an intermediate number of results) can already be efficiently answered with top- search with a small . Thus, in this section we focus on improving queries outside this range, both queries with many results and queries with no results. Intuitively, both types of queries call for some dynamic adjustments to the beam search: for queries with no results, ideally terminating quickly before the end of the search, and for queries with many results, dynamically extending the size of the beam in some way.
4.1. Data Structure and Baseline Algorithm
@@
As the aim of this paper is to investigate whether ANNS graphs can be effectively reused for range search, the data structure used for range queries is an ANNS graph. There are many examples of ANNS graphs in the literature (Subramanya et al., 2019; Fu et al., 2019; Malkov and Yashunin, 2018). For the experiments in this paper we use the in-memory construction of DiskANN, specifically the ParlayANN library (Manohar et al., 2024), but the algorithms we present should work on any single-layer ANNS graph that performs well using the standard beam search, and should be modifiable to multi-layer graphs such as HNSW with little effort.
@@
For our baseline algorithm, due to lack of existing algorithms for range search in high dimensions, we use standard beam search on an ANNS graph. This algorithm is described in detail in many publications, and we also show pseudocode in Algorithm 1.
@@ 4.2. Improving Range Search for Queries with Many Results
We first investigate how to approach graph-based range search for queries that have many results. A standard beam search can only capture as many points as are contained within the beam of size , and thus any queries with significantly more than results cannot achieve high accuracy. The natural question is whether it is possible to distinguish between such queries with more than results and expand the set of returned candidates for those queries, without wasting time on queries with fewer than results. In this section we discuss two algorithmic approaches to this idea.
@@
The first, and perhaps most natural, algorithm extending beam search to queries with variable number results is a dynamic beam search: that is, run the beam search with some starting beam for a certain number of steps, then either terminate or double the beam and continue searching for more candidates. The algorithm terminates if fewer than some fraction of returned candidates are valid range candidates, and continues otherwise. In practice, we found that the value of did not change the results very much, so we set it to as a rule. That algorithm is shown using beam search as a subroutine in Algorithm 5.
@@
Another approach to dynamically resizing the queue length is to run a beam search with unbounded queue length, but much stricter criteria on which nodes are explored. In this approach, candidates are first generated by an initial beam search with beam . If the number of candidates within the radius falls above a certain threshold (again, some fraction of the candidates returned, which we found in practice not to matter significantly), the search moves onto this so-called ”greedy search,” which has an unbounded queue but only adds nodes to the queue if they are within the ball of radius around the query point for the specified radius. This approach takes advantage of the fact that nodes in the same ball of radius are likely to be in connected clusters, and avoids the overhead of the doubling beam search by performing very few distance comparisons to points which are not valid results. See Algorithm 2 for a complete description, and Section 5 for experimental analysis.
@@ 4.3. Improving Range Search for Queries with No Results
This section introduces an algorithm which improves range search for queries with no results. We start from the established paradigm of starting a search with fixed beam and continuing with Algorithm 2 or Algorithm 5 if the returned list contains only candidates within the radius. It would be ideal to terminate the initial beam search early if the query is not finding any results, and the termination condition should use terms that are already computed during the beam search, or inexpensive to compute. A 2020 paper from Li, Zhang, Anderson, and He (Li et al., 2020b) addresses a similar question—finding conditions for terminating top- search early—and finds that several existing terms can be used to predict whether a search can terminate early. Those terms included the coordinates of the query point, the distance from the query point to the current top- neighbor for fixed , and the ratio of the distance from the query point to top- neighbor and the distance from the query point to the start node.
@@
Figure 5(a) histograms the metrics above for MSMARCO-1M at step 20 of a beam search with beam 100. Algorithm 3 shows a natural and general algorithm for imposing an early stopping condition on range search, using any one of these metrics. In all cases, when a search has already found at least one candidate within the specified range, the search does not terminate early. Otherwise, the early stopping condition is applied after a certain number of steps in the beam search. The algorithm terminates when the chosen metric is above a pre-specified cutoff. Thus, in Figure 5(a), we show the same histograms as in Figure 5(b), with points where the search has already found a point inside the radius excluded. The figure also shows vertical lines indicating a potential cutoff for Algorithm 3; points with more than one range result to the right of the line would be “incorrectly” cut off from further search, while points with zero points to the right of the line would be “incorrectly” allowed to proceed instead of terminating early. The goal is to identify a cutoff that yields a more favorable QPS/recall curve. Figure 5(c) shows one choice of metric— on four different datasets.
Based on this data, a larger set of which is shown in Appendix A, we come to three conclusions: first, some datasets show potential for significant improvement using early stopping, while others do not. Second, if a dataset shows potential for improvement using one metric, it also tends to show improvement for all other metrics. Third, using seems to be the best early stopping metric, when examined using the exclusion criteria in Figure 5(b). We hypothesize that is the best metric as it is the most frequently updated and contains the most granular information, but the positive results from other metrics suggests that approaches using machine learning such as those in the Li et al. paper (Li et al., 2020b) might also be useful. We present experimental analysis of this algorithm in Section 5. See Algorithm 4 for an example of how to apply the metric during a search, and Algorithm 6 for an example of the end-to-end algorithm using greedy search.
@@ 5. Experimental Results
In this section, we conduct an experimental evaluation of our algorithms on the nine datasets covered in Section 3. In addition to plotting the QPS and precision tradeoffs for each dataset, we evaluate the time breakdown of each algorithm and draw conclusions about the reasons for their effectiveness.
@@
Our algorithms were implemented as extensions of the ParlayANN (Manohar et al., 2025) library, which is implemented in C++ using ParlayLib (par, [n. d.]) for fork-join parallelism as well as standard building blocks such as sorting and filtering. The code will be made available after the anonymous review period. All experiments were conducted on an Azure Standard_L32s_v3 virtual machine, with a 3rd Generation Intel® Xeon® Platinum 8370C (Ice Lake) processor in a hyper-threaded configuration. It has 32 vCPUs available to the user and a 256 GB main memory. Experiments use all 32 vCPUs unless otherwise stated.
To produce a curve of QPS versus average precision, the starting beamwidth is varied on a number of searches and the Pareto frontier is reported.
@@
We begin our experimental evaluation by plotting queries per second (QPS) versus precision for all nine datasets. Data on all datasets except SSNPP can be found in Figure 6; SSNPP is shown in Figure 7, which illustrates the effects of size scaling on the algorithms using datasets SSNPP and BIGANN. Overall, our algorithms generate improvement over the naive baseline on every dataset, but we find significant diversity in both the magnitude of the improvement and the relative ordering of the two algorithms presented in Section 4. We find that DEEP, BIGANN, and MSTuring experience massive improvement over the baseline (close to 100x improvement in throughput in some cases), with the doubling beam search slightly faster than the greedy search on the 1 million scale, but significantly slower at the 100 million scale. OpenAI also sees large improvement in the realm of one order of magnitude, with greedy search and beam search having virtually identical performance. Wikipedia and MSMarco experience decent improvement with the greedy search somewhat outperforming doubling beam search. GIST achieves such significant improvement over the baseline due to its outlier status with respect to the frequency distribution of its matches, as it is the only dataset with hundreds of points with more than 10000 range results. Greedy search also significantly outperforms doubling beam search for GIST. Text2Image sees modest improvement below 90% recall. SSNPP, which we investigate for three different sizes, first shows little to no improvement at the 1 million scale, then some improvement at the 10 million scale, then significant improvement at size 100 million.
@@
Figure 7 examines the performance of our algorithms on two datasets at three different scales: 1 million, 10 million, and 100 million. Similarly to existing range search benchmarks, we use the same radius for each data scale, so each order of magnitude effectively represents a more dense space, and the average number of results per query point increases. Unsurprisingly, the advantage of our algorithms increases significantly over the baseline, which is only capable of returning up to about 1000 points per query before becoming hopelessly inefficient. Interestingly, we also see that the greedy search seems to have a significant advantage over the doubling beam search as the density of the dataset increases (this is also true of GIST-1M, which has some extraordinarily dense outliers). This is likely because in denser graphs, the greedy search is particularly adept at efficiently finding all points within the radius, which are likely to be in a dense connected cluster in the graph structure. Additional data on size scaling can be found in Appendix A.
@@
To aid in our analysis, we investigate the time spent on different phases of each algorithm in Figure 8. Each algorithm consists of an initial beam search phase—either with or without the early stopping criteria in Algorithm 3—and then a following phase of either greedy search or successive calls to doubling beam search. We show this visual breakdown for three datasets at selected fixed average precision: BIGANN-1M, MSMARCO-1M, and GIST-1M.
The time breakdowns provide significant insight into the results in Figures 6 and 7. First, note that for all datasets, the initial beam search phase is a sizable portion of the overall cost of computation. This is because the initial beam search takes place on all points, but further searches only take place on points with one or more valid range result (since the distances used are exact, a point with zero results will never pass on to the second round of computation).
@@
Next, let us examine the differences between doubling search and greedy search. Doubling search spends a significantly longer portion of time in the second phase of search compared to greedy search, but is capable of achieving higher recall from a smaller starting beam. The characteristics of each individual dataset seem to determine which approach provides the best tradeoff between QPS and average precision. Greedy search requires a larger initial beam to achieve a particular recall value, but the second round of search is extremely cheap as it only visits points within the radius. On the other hand, doubling beam search requires a smaller initial beam, but spends more time in the second phase of search, which performs comparatively more distance comparisons to points which are not valid range results. It also has some overheads related to repeated calls to the beam search function. Based on these factors and the experimental results, it appears that doubling search dominates greedy search when a) a small initial beam is sufficient for doubling search to achieve high recall, and b) a fairly small number of rounds of doubling search are required. These conditions are met with datasets BIGANN-1M (and DEEP-1M, although its time breakdown is not pictured), but GIST-1M requires too many rounds of doubling, and MSMARCO-1M requires too high an initial beam. We also see that in the scaling studies in Figure 7, the greedy search overtakes the doubling search as the density of matches increases, also likely because many rounds of doubling search become necessary.
@@
Finally, we consider the benefit of early stopping. Early stopping clearly decreases the time taken for a fixed beam; for example, in Figure 8(a), early stopping beam search with beam 95 takes less than half the time compared to beam search with beam 95 without early stopping. On the other hand, as illustrated in Figure 8(b), early stopping can make the search significantly less accurate due to points with valid range results being terminated too early. The effects of early stopping on selected datasets are shown in Figure 9 and expanded on in Appendix A; overall, early stopping is beneficial on datasets where significant separation exists between the distributions of points with no results and points with one or more results, and not beneficial otherwise.
@@
We add a final note on comparison with top- search. To investigate this, we ran top-10 searches on OpenAI-1M and MSTuring-1M, and measured the QPS at 90% 10@10 recall. For OpenAI-1M, we found the QPS at 90% recall was around 5000, compared to around 10000 QPS for range search at 90% precision. For MSTuring-1M, the QPS at 90% recall for top-10 search was about 30000, compared to over 100000 for range search using our algorithms. This suggests that range benchmarking may actually be thought of as an easier problem than top- search.
@@ Acknowledgements.
- Douze et al. (2024) Matthijs Douze, Alexandr Guzhva, Chengqi Deng, Jeff Johnson, Gergely Szilvasy, Pierre-Emmanuel Mazaré, Maria Lomeli, Lucas Hosseini, and Hervé Jégou. 2024. The Faiss library. CoRR abs/2401.08281 (2024). https://doi.org/10.48550/ARXIV.2401.08281 arXiv:2401.08281
- Douze et al. (2021) Matthijs Douze, Giorgos Tolias, Ed Pizzi, Zoë Papakipos, Lowik Chanussot, Filip Radenovic, Tomás Jenícek, Maxim Maximov, Laura Leal-Taixé, Ismail Elezi, Ondrej Chum, and Cristian Canton-Ferrer. 2021. The 2021 Image Similarity Dataset and Challenge. CoRR abs/2106.09672 (2021). arXiv:2106.09672 https://arxiv.org/abs/2106.09672
- Fu et al. (2019) Cong Fu, Chao Xiang, Changxu Wang, and Deng Cai. 2019. Fast Approximate Nearest Neighbor Search With The Navigating Spreading-out Graph. Proc. VLDB Endow. 12, 5 (2019), 461–474. https://doi.org/10.14778/3303753.3303754
@@
- Indyk and Motwani (1998) Piotr Indyk and Rajeev Motwani. 1998. Approximate Nearest Neighbors: Towards Removing the Curse of Dimensionality. In Proceedings of the Thirtieth Annual ACM Symposium on the Theory of Computing, Dallas, Texas, USA, May 23-26, 1998, Jeffrey Scott Vitter (Ed.). ACM, 604–613. https://doi.org/10.1145/276698.276876
@@
- Li et al. (2020b) Conglong Li, Minjia Zhang, David G. Andersen, and Yuxiong He. 2020b. Improving Approximate Nearest Neighbor Search through Learned Adaptive Early Termination. In Proceedings of the 2020 International Conference on Management of Data, SIGMOD Conference 2020, online conference [Portland, OR, USA], June 14-19, 2020, David Maier, Rachel Pottinger, AnHai Doan, Wang-Chiew Tan, Abdussalam Alawini, and Hung Q. Ngo (Eds.). ACM, 2539–2554. https://doi.org/10.1145/3318464.3380600
@@
- Manohar et al. (2024) Magdalen Dobson Manohar, Zheqi Shen, Guy E. Blelloch, Laxman Dhulipala, Yan Gu, Harsha Vardhan Simhadri, and Yihan Sun. 2024. ParlayANN: Scalable and Deterministic Parallel Graph-Based Approximate Nearest Neighbor Search Algorithms. In Proceedings of the 29th ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming, PPoPP 2024, Edinburgh, United Kingdom, March 2-6, 2024, Michel Steuwer, I-Ting Angelina Lee, and Milind Chabbi (Eds.). ACM, 270–285. https://doi.org/10.1145/3627535.3638475
@@
- Simhadri et al. (2021a) Harsha Vardhan Simhadri, George Williams, Martin Aumüller, Matthijs Douze, Artem Babenko, Dmitry Baranchuk, Qi Chen, Lucas Hosseini, Ravishankar Krishnaswamy, Gopal Srinivasa, Suhas Jayaram Subramanya, and Jingdong Wang. 2021a. Results of the NeurIPS’21 Challenge on Billion-Scale Approximate Nearest Neighbor Search. In NeurIPS 2021 Competitions and Demonstrations Track, 6-14 December 2021, Online (Proceedings of Machine Learning Research, Vol. 176), Douwe Kiela, Marco Ciccone, and Barbara Caputo (Eds.). PMLR, 177–189. https://proceedings.mlr.press/v176/simhadri22a.html
- Subramanya et al. (2019) Suhas Jayaram Subramanya, Devvrit, Harsha Vardhan Simhadri, Ravishankar Krishnaswamy, and Rohan Kadekodi. 2019. Rand-NSG: Fast Accurate Billion-point Nearest Neighbor Search on a Single Node. In Advances in Neural Information Processing Systems 32: Annual Conference on Neural Information Processing Systems 2019, NeurIPS 2019, December 8-14, 2019, Vancouver, BC, Canada, Hanna M. Wallach, Hugo Larochelle, Alina Beygelzimer, Florence d’Alché-Buc, Emily B. Fox, and Roman Garnett (Eds.). 13748–13758. https://proceedings.neurips.cc/paper/2019/hash/09853c7fb1d3f8ee67a61b6bf4a7f8e6-Abstract.html
- Szilvasy et al. (2024) Gergely Szilvasy, Pierre-Emmanuel Mazaré, and Matthijs Douze. 2024. Vector search with small radiuses. CoRR abs/2403.10746 (2024). https://doi.org/10.48550/ARXIV.2403.10746 arXiv:2403.10746
- Wang et al. (2013) Jie Wang, Jian Lu, Zheng Fang, Tingjian Ge, and Cindy X. Chen. 2013. PL-Tree: An Efficient Indexing Method for High-Dimensional Data. In Advances in Spatial and Temporal Databases - 13th International Symposium, SSTD 2013, Munich, Germany, August 21-23, 2013. Proceedings (Lecture Notes in Computer Science, Vol. 8098), Mario A. Nascimento, Timos K. Sellis, Reynold Cheng, Jörg Sander, Yu Zheng, Hans-Peter Kriegel, Matthias Renz, and Christian Sengstock (Eds.). Springer, 183–200. https://doi.org/10.1007/978-3-642-40235-7_11
@@
- Wang et al. (2021) Mengzhao Wang, Xiaoliang Xu, Qiang Yue, and Yuxiang Wang. 2021. A Comprehensive Survey and Experimental Comparison of Graph-Based Approximate Nearest Neighbor Search. Proc. VLDB Endow. 14, 11 (2021), 1964–1978. https://doi.org/10.14778/3476249.3476255

### v3

@@ 4. Range Search on Graph-Based ANNS Indices
This section introduces our range search algorithms, starting with the data structure and baseline beam search. Since beam search performs well for top- queries, some range queries with an intermediate number of results can be efficiently handled using a small . Thus, we focus on improving queries at the extremes—those with many results and those with none. Intuitively, both require dynamic adjustments: terminating early for queries with no results and dynamically extending the beam for those with many.
4.1. Data Structure and Baseline Algorithms
@@
As the aim of this paper is to investigate whether ANNS graphs can be effectively reused for range search, the data structure used for range queries is an ANNS graph. There are many examples of ANNS graphs in the literature (Subramanya et al., 2019; Malkov and Yashunin, 2018). For the experiments in this paper we use the in-memory construction of DiskANN, specifically the ParlayANN library (Manohar et al., 2024), but the algorithms we present should work on any single-layer ANNS graph that uses the standard beam search algorithm, and should be modifiable to multi-layer graphs such as HNSW with little effort.
@@
For our baseline algorithms, due to lack of existing algorithms for range search on graph-based indices, we use standard beam search on an ANNS graph as one baseline. This algorithm is described in detail in many publications (Manohar et al., 2024; Subramanya et al., 2019), and we also show pseudocode in Algorithm 1. We also build on the doubling search baseline suggested in the Starling system (Wang et al., 2024), re-implementing it within the ParlayANN library for the in-memory setting for the fairest possible comparison.
@@ 4.2. Improving Range Search for Queries with Many Results
We first investigate how to approach graph-based range search for queries that have many results. A standard beam search can only capture as many points as are contained within the beam of size , and thus any queries with significantly more than results cannot achieve high accuracy. The natural question is whether it is possible to identify such queries with more than results and expand the set of returned candidates for only those queries. In this section we discuss two algorithmic approaches to this idea.
@@
The first, and perhaps most natural, algorithm extending beam search to queries with variable number results is a dynamic beam search: that is, run the beam search with some starting beam , then either terminate or double the beam and continue searching for more candidates. This algorithm was first proposed in the paper introducing the Starling system (Wang et al., 2024). The algorithm terminates if fewer than some fraction of returned candidates are valid range candidates, and continues otherwise. In practice, we found that the value of did not change the results very much, so we set it to as a rule (this finding seems to differ from that of the Starling authors, which may be attributed either to implementation distances or the in-memory setting). That algorithm is shown using beam search as a subroutine in Algorithm 3. In addition to rewriting the doubling search within the in-memory ParlayANN library, we made some minor optimizations such as using a faster approximate hash table for visited set lookups. For the fairest possible comparison, we also augment it with the early stopping heuristic introduced in Section 4.3 during our experimental evaluation, but we note that the version in their paper contains no such heuristic.
@@
Another approach to dynamically resizing the queue length is to run a beam search with unbounded queue length, but much stricter criteria on which nodes are explored. Candidates are first generated by an initial beam search with beam . If the number of candidates within the radius falls above a certain threshold (again, some fraction of the candidates returned), we next call our so-called greedy search routine, which has an unbounded queue but only adds nodes to the queue if they are within the ball of radius around the query point for the specified radius. This approach takes advantage of the fact that nodes in the same ball of radius are likely to be in connected clusters, and avoids the overhead of the doubling beam search by performing very few distance comparisons to points which are not valid results. See Algorithm 2 in the supplementary material for a complete description, and Section 5 for experimental analysis.
@@ 4.3. Improving Range Search for Queries with No Results
This section introduces an algorithm to improve range search for queries with no results. The search begins with a fixed beam , continuing with Algorithm 2 or Algorithm 3 if only candidates within the radius are returned. Ideally, the initial beam search should terminate early when no results are found, using terms that are inexpensive to compute. A 2020 paper by Li et al. (Li et al., 2020b) explores similar termination criteria for top- search, identifying several predictive terms: query point coordinates, distance to the current top- neighbor for fixed , and the ratio of that distance to the query point’s distance from the start node.
@@
A natural and general approach to an early stopping condition for range search, using a variety of metrics, is to apply the early stopping metric after a certain number of steps in the beam search. The search does not terminate early if it has already found at least one candidate in the specified range. Figure 3(c) shows one choice of metric— on four different datasets. Algorithm 4 shows formally how this condition can be applied.
Based on the histogram data of early stopping conditions, a larger set of which is shown in the supplemental material, we come to three conclusions: first, some datasets show potential for significant improvement using early stopping, while others do not. Second, if a dataset shows potential for improvement using one metric, it also tends to show improvement for all other metrics. Third, using seems to be the best early stopping metric. We hypothesize that is the best metric as it is the most frequently updated and contains the most granular information, but the positive results from other metrics suggests that approaches using machine learning such as those in the Li et al. paper (Li et al., 2020b) might also be useful. We present experimental analysis of this algorithm in Section 5. See Algorithm 5 for an example of how to apply the metric during a search, and Algorithm 6 for an example of the end-to-end algorithm using greedy search.
@@ 5. Experimental Results
We evaluate our algorithms on the nine datasets covered in Section 3.
@@
The algorithms introduced in Section 4 extend the ParlayANN (Manohar et al., 2025) library, implemented in C++ with ParlayLib (par, [n. d.]) for fork-join parallelism and standard parallel primitives such as sorting and filtering. ParlayANN features a top-ranked DiskANN implementation, currently among the top five on ANN Benchmarks (Aumüller et al., 2021). The code is publicly available at https://github.com/cmuparlay/ParlayANN/tree/rangetest2. Most experiments ran on an Azure Standard_L32s_v3 VM with a 3rd Gen Intel® Xeon® Platinum 8370C processor, 32 vCPUs, and 256 GB memory, utilizing all vCPUs unless noted. Billion-scale tests were conducted on a 72-core Dell R930 with 4x Intel® Xeon® E7-8867 v4 processors, each with 18 cores, 2.4GHz, and 45MB L3 cache, and 1 TB main memory.
To produce a curve of QPS versus average precision, the starting beamwidth is varied on a number of searches and the Pareto frontier is reported. The build parameters used for the DiskANN graphs, as well as the early stopping threshold for each dataset, are reported in the supplemental material. We use scalar quantization to 8 bits for all floating-point datasets, with re-ranking to ensure that all reported results are within the radius.
Competitors
To evaluate our algorithms against another standard baseline, we also use the range search algorithms from the well-known FAISS-IVF library (Douze et al., 2024). At a high level, FAISS-IVF uses an inverted file index produced via clustering, with the option to construct an HNSW index over the cluster centroids to help select clusters more efficiently at query time. A query probes a selected number of cells and returns the candidates within the radius. The parameters chosen for FAISS, along with information on how we configured them, are shown in the supplemental material. To produce a curve of QPS versus average precision for FAISS, we varied both the number of cells probed and the of the centroid graph and report the Pareto frontier.
@@
Our experimental evaluation plots queries per second (QPS) versus precision for nine datasets (Figure 4). Figure 5 examines size scaling effects for SSNPP and BIGANN, from 10 million to 1 billion. While our algorithms improve upon the naive baseline across all datasets, the degree of improvement and algorithm rankings vary. DEEP, BIGANN, and MSTuring achieve near-100x throughput gains over FAISS-IVF and the beam search baseline, with doubling beam search similar than greedy search at 1 million scale but up to 10x slower on larger datasets. OpenAI sees a 10x improvement over beam search and FAISS, with both doubling and greedy search performing similarly. Wikipedia and MSMARCO-WebSearch improve moderately over the FAISS and beam search baselines, with greedy search slightly ahead of doubling. GIST stands out due to its unique frequency distribution, as the only dataset with hundreds of points exceeding 10,000 range results, making greedy search more effective. Text2Image shows modest gains below 90% recall. SSNPP, analyzed across four scales, exhibits little improvement at 1 million but increasingly benefits from the specialized range search algorithms as dataset size grows.
Comparison with FAISS-IVF
Greedy search and doubling search are overall faster than FAISS-IVF, with occasional exceptions at the highest recall values. Interestingly, FAISS-IVF is faster than the beam search baseline about half the time. These results reflect the fact that the FAISS algorithm is more intrinsically suited to range search than the beam search baseline, since it considers every point in each cell as a potential range candidate. However, greedy search and doubling search with early stopping are overall more successful at precisely searching the correct number of points for each range query.
@@
Figure 5 examines the performance of our algorithms on two datasets at three different scales: 10 million, 100 million, and 1 billion. Similarly to existing range search benchmarks, we use the same radius for each data scale, so each order of magnitude effectively represents a more dense space, and the average number of results per query point increases. Unsurprisingly, the advantage of greedy search and doubling search increase significantly over the beam search baseline, which is only capable of returning up to about 1000 points per query before becoming hopelessly inefficient. Interestingly, we also see that the greedy search seems to have the most significant advantage over the doubling beam search as the density of the dataset increases (this is also true of GIST-1M, which has some extraordinarily dense outliers). This is likely because in denser graphs, the greedy search is particularly adept at efficiently finding all points within the radius, which are likely to be in a dense connected cluster in the graph structure. In the supplemental material we investigate this hypothesis directly by fixing one dataset and measuring performance of both algorithms while varying the chosen radius. Greedy search had better performance for larger choices of radius (indicating more matches per point on average).
@@
To aid in our analysis, we investigate the time spent on different phases of each algorithm in Figure 6. Each algorithm consists of an initial beam search phase—either with or without the early stopping criteria in Algorithm 4—and then a following phase of either greedy search or successive calls to doubling beam search. We show this visual breakdown for three datasets at selected fixed average precision: BIGANN-1M, BIGANN-100M, and MSMarcoWebSearch-1M. The time breakdowns provide significant insight into the results in Figures 4 and 5. First, note that for all datasets, the initial beam search phase is a sizable portion of the overall cost of computation. This is because the initial beam search takes place on all points, but further searches only take place on points with one or more valid range result (since the distances used are exact, a point with zero results will never pass on to the second round of computation).
@@
Next, we examine the differences between doubling and greedy search. Doubling search spends more time in the second phase but achieves higher recall from a smaller starting beam. Dataset characteristics determine the best tradeoff between QPS and average precision. Greedy search requires a larger initial beam but has a cheaper second round since it only visits points within the radius. In contrast, doubling search starts with a smaller beam but spends more time in the second phase, performing additional distance comparisons and incurring overhead from repeated beam search calls. Experimental results suggest doubling search dominates greedy search when (a) a small initial beam is enough for doubling search to achieve high recall, and (b) few rounds of doubling are needed. These conditions hold for BIGANN-1M (and DEEP-1M), but GIST-1M requires excessive doubling rounds, and MSMARCOWebSearch-1M demands a large initial beam. Scaling studies in Figure 5 show greedy search overtakes doubling search as match density increases, likely due to the growing number of doubling rounds required.
Effects of Increasing Radius
To better isolate the reason for greedy search’s superior performance, in Figure 7, we measure the effect of varying the radius for a fixed dataset for three selected values of average precision, annotating each data point with the beam used for the initial beam search. To aid the reader’s understanding, instead of marking the x-axis with the raw value of the radius used, we instead measure the average number of matches per query induced by the chosen radius. We see that in most cases greedy search outperforms or performs equally to doubling search, with the effect becoming more pronounced as the number of matches per point increases. This confirms the observations in the size scaling study.
@@
Next, we examine the impact of early stopping. It significantly reduces search time for a fixed beam; for example, in Figure 6(b), early stopping with beam 44 takes only about twice the time as a search with beam 2 and no early stopping. However, as Figure 6(c) shows, early stopping can reduce accuracy by prematurely terminating points with valid range results. Appendix A details its effects on selected datasets. Overall, early stopping is beneficial when there is clear separation between points with no results and those with one or more results.
@@
We add a final note on comparison with top- search. To investigate this, we ran top-10 searches on OpenAI-1M and MSTuring-1M, and measured the QPS at 90% recall. For OpenAI-1M, the QPS at 90% recall was around 5000, compared to around 10000 QPS for range search at 90% precision. For MSTuring-1M, the QPS at 90% recall for top-10 search was about 30000, compared to over 100000 for range search using our algorithms. This suggests range search may actually be an easier problem than top- search.
@@ 6. Conclusion
Finally, there is a notable lack of benchmarks designed for range searching. Contributions of new datasets with natural range search benchmarks would significantly further this area of exploration.
@@ Acknowledgements.
- Aumüller et al. (2021) Martin Aumüller, Erik Bernhardsson, and Alexander Faithfull. 2021. ANN Benchmarks. http://ann-benchmarks.com/.
@@
- Douze (2022) Matthijs Douze. 2022. Faiss Wiki. Webpage. Retrieved April 29, 2025 from https://github.com/facebookresearch/faiss/wiki
- Douze et al. (2024) Matthijs Douze, Alexandr Guzhva, Chengqi Deng, Jeff Johnson, Gergely Szilvasy, Pierre-Emmanuel Mazaré, Maria Lomeli, Lucas Hosseini, and Hervé Jégou. 2024. The Faiss library. CoRR (2024). https://doi.org/10.48550/arXiv.2401.08281
- Douze et al. (2021) Matthijs Douze, Giorgos Tolias, Ed Pizzi, Zoë Papakipos, Lowik Chanussot, Filip Radenovic, Tomás Jenícek, Maxim Maximov, Laura Leal-Taixé, Ismail Elezi, Ondrej Chum, and Cristian Canton-Ferrer. 2021. The 2021 Image Similarity Dataset and Challenge. CoRR (2021). https://arxiv.org/abs/2106.09672
- Engels et al. (2024) Joshua Engels, Benjamin Landrum, Shangdi Yu, Laxman Dhulipala, and Julian Shun. 2024. Approximate Nearest Neighbor Search with Window Filters. In Forty-first International Conference on Machine Learning, ICML 2024, Vienna, Austria, July 21-27, 2024. OpenReview.net. https://openreview.net/forum?id=8t8zBaGFar
- Fu et al. (2019) Cong Fu, Chao Xiang, Changxu Wang, and Deng Cai. 2019. Fast Approximate Nearest Neighbor Search With The Navigating Spreading-out Graph. Proc. VLDB Endow. 12, 5 (2019), 461–474. http://www.vldb.org/pvldb/vol12/p461-fu.pdf
@@
- Indyk and Motwani (1998) Piotr Indyk and Rajeev Motwani. 1998. Approximate Nearest Neighbors: Towards Removing the Curse of Dimensionality. In Proceedings of the Thirtieth Annual ACM Symposium on the Theory of Computing, Dallas, Texas, USA, May 23-26, 1998. ACM, 604–613. https://doi.org/10.1145/276698.276876
@@
- Li et al. (2020b) Conglong Li, Minjia Zhang, David G. Andersen, and Yuxiong He. 2020b. Improving Approximate Nearest Neighbor Search through Learned Adaptive Early Termination. In Proceedings of the 2020 International Conference on Management of Data, SIGMOD Conference 2020, online conference [Portland, OR, USA], June 14-19, 2020. ACM, 2539–2554. https://doi.org/10.1145/3318464.3380600
@@
- Manohar et al. (2024) Magdalen Dobson Manohar, Zheqi Shen, Guy E. Blelloch, Laxman Dhulipala, Yan Gu, Harsha Vardhan Simhadri, and Yihan Sun. 2024. ParlayANN: Scalable and Deterministic Parallel Graph-Based Approximate Nearest Neighbor Search Algorithms. In Proceedings of the 29th ACM SIGPLAN Annual Symposium on Principles and Practice of Parallel Programming, PPoPP 2024, Edinburgh, United Kingdom, March 2-6, 2024. ACM, 270–285. https://doi.org/10.1145/3627535.3638475
@@
- Simhadri et al. (2021a) Harsha Vardhan Simhadri, George Williams, Martin Aumüller, Matthijs Douze, Artem Babenko, Dmitry Baranchuk, Qi Chen, Lucas Hosseini, Ravishankar Krishnaswamy, Gopal Srinivasa, Suhas Jayaram Subramanya, and Jingdong Wang. 2021a. Results of the NeurIPS’21 Challenge on Billion-Scale Approximate Nearest Neighbor Search. In NeurIPS 2021 Competitions and Demonstrations Track, 6-14 December 2021, Online (Proceedings of Machine Learning Research, Vol. 176). PMLR, 177–189. https://proceedings.mlr.press/v176/simhadri22a.html
- Subramanya et al. (2019) Suhas Jayaram Subramanya, Devvrit, Harsha Vardhan Simhadri, Ravishankar Krishnaswamy, and Rohan Kadekodi. 2019. Rand-NSG: Fast Accurate Billion-point Nearest Neighbor Search on a Single Node. In Advances in Neural Information Processing Systems 32: Annual Conference on Neural Information Processing Systems 2019, NeurIPS 2019, December 8-14, 2019, Vancouver, BC, Canada. 13748–13758. https://proceedings.neurips.cc/paper/2019/hash/09853c7fb1d3f8ee67a61b6bf4a7f8e6-Abstract.html
- Szilvasy et al. (2024) Gergely Szilvasy, Pierre-Emmanuel Mazaré, and Matthijs Douze. 2024. Vector search with small radiuses. CoRR abs/2403.10746 (2024). https://doi.org/10.48550/arXiv.2403.10746
- Wang et al. (2013) Jie Wang, Jian Lu, Zheng Fang, Tingjian Ge, and Cindy X. Chen. 2013. PL-Tree: An Efficient Indexing Method for High-Dimensional Data. In Advances in Spatial and Temporal Databases - 13th International Symposium, SSTD 2013, Munich, Germany, August 21-23, 2013. Proceedings (Lecture Notes in Computer Science, Vol. 8098). Springer, 183–200. https://doi.org/10.1007/978-3-642-40235-7_11
@@
- Wang et al. (2021) Mengzhao Wang, Xiaoliang Xu, Qiang Yue, and Yuxiang Wang. 2021. A Comprehensive Survey and Experimental Comparison of Graph-Based Approximate Nearest Neighbor Search. Proc. VLDB Endow. 14, 11 (2021), 1964–1978. http://www.vldb.org/pvldb/vol14/p1964-wang.pdf